mhchen:

I hope question cove is smart enough to answer this: Why is the domain of tan( arcsin(x/2) ) equal to (-2,2) I know that the domain of arcsin(x/2) is (-2,2), I don't know why the tangent function didn't change it.

AMQ:

a few users are, but the aren't online at the moment :/

mhchen:

sigh....please don't make me go back to the other dumb sites

AMQ:

like what? brainly and pa? lol

Ultrilliam:

Uhm, if mathmale or sat where on right now, they could help :(

AMQ:

*were also, @tranquility could possibly answer it :/

Ultrilliam:

Wait, I have someone who might be able to help

mhchen:

Mathmale gives me PTSD

Ultrilliam:

I'm talking to someone who is guiding me through how to answer this for you :)

Ultrilliam:

The tangent function does change the domain. for the arcsin(x/2), the domain includes points -2 and 2. while when you include the tangent, those x=-2 and x=2 areas/lines become asymptotes and at no longer included

mhchen:

Oh.........RIGHT yes

mhchen:

but ask your boyfriend why the tangent doesn't change the (-2,2) into anything else, like why is it still 2.

Ultrilliam:

First of it's a girl, and she is way older then me. and second, i'm asking her now.

Ultrilliam:

Sorry for the long wait lol

Ultrilliam:

it's not that it doesn't change the domain or range - the idea behind it does change the arcsin(x/2)'s domain of [-2,2] because of how the stuff in the parenthesis of arcsin is limited from -1 to 1 meanwhile, tan = opp/adj from arcsin(x/2) we get that opp=x and hyp=2 so from that we can get that adj=sqrt(4-x^2) so the tan(arcsin(x/2))=x/(sqrt(4-x^2)) and that is where that domain comes from

celticcat:

Yes - there are restrictions on the domain of the tangent because of the restrictions on the sine

mhchen:

I LOVE YOU ULTRILLIAM AND YOUR BOYFRIEND

mhchen:

I have more hope for qc now

Ultrilliam:

it still isn't my boyfriend :| once again, its a girl. GIRL And i'm glad! :D Welcome aboard! ^_^

Elsa213:

Is it a girlfriend? o;

Elsa213:

That's why Aeon was created. cx

Ultrilliam:

Oh boy, not at all Elsa lol